OpenStudy (anonymous):

The formula for the perimeter of a rectangle is P = 2l + 2w. If P = 104 and l = 13, what is the length of w? A. 30 B. 33 C. 36 D. 39

OpenStudy (anonymous):

What part of it you need help with?

OpenStudy (arnavguddu):

P=2(l+w) or P/2= l+w. so w = P/2-l now put in the values of p and l in above u will have the answer

OpenStudy (anonymous):

thanks :D